OpenStudy (anonymous):

NEED HELP ASAP!!! WILL FAN MEDAL Which equation shows the substitution method being used to solve the system of linear equations? x – y = 7 x = y + 3 A. x – (y + 3) = 7 B. (y + 3) – y = 7 C. x = (7 – x) + 3 D. x – y = y + 3

OpenStudy (immanuelv):

which one do you think it is?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

x = y + 3 plug in that for the 1 equation given for x

OpenStudy (anonymous):

HUH

OpenStudy (anonymous):

im so lost

OpenStudy (immanuelv):

just follow what @RyleeC said. You need to plug in y +3 to the x in the first equation that will lead you to your answer

OpenStudy (anonymous):

@fateal @31356 @Secret-Ninja @myininaya

OpenStudy (anonymous):

ok still lost

OpenStudy (immanuelv):

at this point idk how to answer without giving the answer :S

OpenStudy (anonymous):

thats fine i need a better understanding any way so i just need a better explanation

OpenStudy (immanuelv):

let me try again Equation 1) x – y = 7 Equation 2) x = y + 3 Theses are the two equations right? substitution method can be used like this you plug in equation 2 into equation 1 so in equation 2 you have x = y +3 so you can just plug in equation 2 into equation 1 so it will be x - y = 7 here x is going to get substituted (y+3) - y = 7

OpenStudy (anonymous):

k thx, let me work it out and can you check my answer

OpenStudy (anonymous):

b or c

OpenStudy (immanuelv):

you got this. Hope my explanation didn't complicate it for you

OpenStudy (anonymous):

it didnt but is it b or c?

OpenStudy (immanuelv):

answer is B

OpenStudy (anonymous):

ok thx thx thx soo much